# Quantum Convolutional Neural Networks for Jet Images Classification

### v2

@@ III.3 DEA circuit
For this task, we select the setup that resulted in the best accuracies for both QCNNs and CNNs. This setup, as shown in Table 1, corresponds to the use of MSE loss along with the HEE1 encoding. After performing a DEA check on the $SU(4)$ QCNN circuit with originally 48 parameters, 17 of them were found to be redundant, leaving a total of 31 parameters in the optimized circuit. This is compatible with the dimension of the state space ($12^{Q+1}-$ 1, where $Q$ is the number of qubits in the system). Therefore, the circuit is minimal in the set of trainable parameters and maximally expressive.
Test accuracy results of the MSE loss averaged over 1000 runs for $SU(4)$ and $SO(4)$ QCNN circuits before and after applying DEA, compared to CNN accuracy can be found in Table 3 and Fig. 8. The accuracy values at epoch 30 in Table 3 indicate a clear difference between the case when we have comparable number of parameters (31) of the DEA circuit with the CNN model (33 parameters) where the DEA circuit has higher accuracy value. The curves in Fig 8 also depict this result between all QCNN circuits (with and without DEA) compared to the CNN accuracy. We may notice slight differences between the $SU(4)$ circuit with and without DEA and the $SO(4)$ circuit at early epochs. The $SU(4)$ circuit without DEA has more parameters (48) which help it converge to the solution after fewer training epochs compared to the case with less parameters in the circuit with DEA (31). Both $SO(4)$ circuits, with and without DEA have a similar performance, and again we notice the slightly higher accuracy at early epochs for the circuit with the higher number of parameters. However, the $SU(4)$ circuit with DEA, has much less parameters compared to its original circuit and yet it shows comparable accuracy values. The $SO(4)$ circuit has 30 parameters, which is very close to the 31 parameters in the $SU(4)$ circuit with DEA. Notably, the DEA circuit shows comparable performances and even improved accuracy in the early epochs, suggesting that DEA can reduce the number of trainable parameters while maintaining strong performance.
[…]
$SU(4)$ without DEA 48 99.93 ± 0.01
$SU(4)$ with DEA 31 99.86 $\pm$ 0.01
$SO(4)$ without DEA 30 99.87 ± 0.01
$SO(4)$ with DEA 25 99.85 ± 0.01
CNN 33 93.08 ± 0.55
@@
IV Conclusion and outlook
In this paper, we utilized two QCNN architectures one with $SO(4)$ convolutional circuit and the other built with an $SU(4)$ convolutional circuit. We varied the setups of each architecture to test different loss functions, encodings, and batch-sizes. We compared each setup with a corresponding CNN model. From the preliminary results we got for our quantum and classical models, we found that the QCNN, in the absence of noise, performs better than the corresponding CNN model for specific choices of data encoding, batch-size, loss function, and convolutional block. However, it is important to emphasize that the system we studied here was relatively simple, both in QCNN and CNN. As a result, it was possible to train the classical model with a very small number of parameters, whereas the usual case of training CNN models is to use a high number of parameters that may reach up to millions. Besides, we only performed the noiseless simulation and the effects from noise would spoil the QCNN performance unless we can tame them.

### v3

@@ IV.3 DEA circuit
For this task, we select the setup that resulted in the best accuracies for both QCNNs and CNNs. This setup, as shown in Table 1, corresponds to the use of MSE loss along with the HEE1 encoding. After performing a DEA check on the $SU(4)$ QCNN circuit with originally 48 parameters, 17 of them were found to be redundant, leaving a total of 31 parameters in the optimized circuit. This is compatible with the dimension of the state space ($2^{Q+1}-1$, where $Q$ is the number of qubits in the system). Therefore, the circuit is minimal in the set of trainable parameters and maximally expressive.
Test accuracy results of the MSE loss averaged over 1000 runs for $SU(4)$ and $SO(4)$ QCNN circuits before and after applying DEA, compared to CNN accuracy can be found in Table 3 and Fig. 13. The accuracy values at epoch 30 in Table 3 indicate a clear difference between the case when we have comparable number of parameters (31) of the DEA circuit with the CNN model (33 parameters) where the DEA circuit has higher accuracy value. The curves in Fig. 13 also depict this result between all QCNN circuits (with and without DEA) compared to the CNN accuracy. We may notice slight differences between the $SU(4)$ circuit with and without DEA and the $SO(4)$ circuit at early epochs. The $SU(4)$ circuit without DEA has more parameters (48) which help it converge to the solution after fewer training epochs compared to the case with less parameters in the circuit with DEA (31). Both $SO(4)$ circuits, with and without DEA have a similar performance, and again we notice the slightly higher accuracy at early epochs for the circuit with the higher number of parameters. However, the $SU(4)$ circuit with DEA, has much less parameters compared to its original circuit and yet it shows comparable accuracy values. The $SO(4)$ circuit has 30 parameters, which is very close to the 31 parameters in the $SU(4)$ circuit with DEA. Notably, the DEA circuit shows comparable performances and even improved accuracy in the early epochs, suggesting that DEA can reduce the number of trainable parameters while maintaining strong performance.
Circuit structure params. Acc ($\%$)
$SU(4)$ without DEA 48 99.95 ± 0.01
$SU(4)$ with DEA 31 99.89 $\pm$ 0.01
$SO(4)$ without DEA 30 99.95 ± 0.01
$SO(4)$ with DEA 25 99.93 ± 0.01
CNN 33 93.08 ± 0.44
@@
V Conclusion and outlook
In this paper, we utilized two QCNN architectures one with $SO(4)$ convolutional circuit and the other built with an $SU(4)$ convolutional circuit. We varied the setups of each architecture to test different loss functions, encodings, and batch-sizes. We compared each setup with a corresponding CNN model. From the preliminary results we got for our quantum and classical models, we found that the QCNN, in the absence of noise, performs better than the corresponding CNN model for most choices of data encoding, batch-size, loss function, and convolutional block involved. However, it is important to emphasize that the system we studied here was relatively simple, both in QCNN and CNN. As a result, it was possible to train the classical model with a very small number of parameters, whereas the usual case of training CNN models is to use a high number of parameters that may reach up to millions. Besides, we only performed the noiseless simulation and the effects from noise would spoil the QCNN performance unless we can tame them.
@@ V Conclusion and outlook
It is worth noting that, despite the improved accuracy we found for QCNN over the CNN model, we are not claiming that our model outperforms all the existing models that did not achieve comparable accuracies for the top-tagging task in general. This is because the accuracy results depend heavily on the specific dataset used and the details of its preprocessing. Hence, a direct, one-to-one comparison is not possible unless fair training conditions with the same dataset was performed. That said, we believe our model holds promising potential for jet image classification tasks in a more general manner. In particular, if the dataset was pre-processed such that when PCA is applied, the principal components can still capture most of the important features existed in the image.
@@
Code availability
The code used for data preprocessing, model implementation, and training, which produced the results reported in this work, is available at https://github.com/haelhag/QCNNs4JetImagesClassification.git. The repository includes scripts and documentation to allow reproduction of the results and further investigation.